# A virtual Clebsch-Gordan formula for $\operatorname{SL}(2,\mathbb{Z})$

### v5

@@ 3. On the ”complete reduciblilty” property for $\operatorname{SL}(2,\mathbb{Z})$
where $g$ is the canonical multiplication map and where $f(v):=zv$ where $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}$. It follows the sequence is an exact sequence of $\operatorname{SL}(2,A)$-comodules where $A$ is an arbitrary commutative unital ring.
@@
and it follows $a:=\frac{n}{n+1}\in A$. From this it follows $\frac{1}{n+1}\in A$. Hence the projection map $\pi$ does not have a section $s$ if $A=\mathbb{Z}$. Hence the group scheme $\operatorname{SL}(2,\mathbb{Z})$ does not satisfy the ”complete reducibility” property. If $(p)\subseteq\mathbb{Z}$ is a prime ideal and if $A:=\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at $(p)$, it follows the sequence
@@
does not split since $p$ is not a unit in $A$. Hence $\operatorname{SL}(2,\mathbb{Z}_{(p)})$ does not satisfy the ”complete reducibility” property. If $G:=\operatorname{SL}(2,\mathbb{Z}[\frac{1}{m}])$ there is always an integer $n$ where the projection map $\pi_{n}$ does not split, hence $G$ does not satisfy the ”complete reducibility” property.
@@
Let $m$ be any positive integer and let $p\in\mathbb{Z}$ be a prime number. The group schemes $\operatorname{SL}(2,\mathbb{Z}),\operatorname{SL}(2,\mathbb{Z}_{(p)})$ and $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{m}])$ do not satifsy the complete reducibility property.
@@
If $k$ is a field of characteristic zero and if $V$ is any finite dimensional $\operatorname{SL}(2,k)$-module, it follows there is a direct sum decomposition $V\cong\oplus_{i\in I}\operatorname{S}^{d_{i}}(V)$ of representations. The symmetric powers $\operatorname{S}^{d}(V)$ are irreducible when $V:=k\{e_{1},e_{2}\}$ is the standard $\operatorname{SL}(2,k)$-module. Proposition 3.7 proves this property does not hold for the group scheme $\operatorname{SL}(2,\mathbb{Z})$. The symmetric power of the standard module $\operatorname{S}^{d}(V)$ is not irreducible and there is no decomposition of the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ into a direct sum of irreducible modules. By the above example it follows $V\otimes\operatorname{S}^{n}(V)$ is not isomorphic to the direct sum $\operatorname{S}^{n-1}(V)\oplus\operatorname{S}^{n+1}(V)$ as comodules. Hence the ”classification” of $\operatorname{SL}(2,\mathbb{Z})$-modules of finite rank over $\mathbb{Z}$ is more complicated than the case of $\operatorname{SL}(2,k)$.
@@
Let $V$ be the standard comodule on $\operatorname{SL}(2,\mathbb{Z})$ from Definition 2.1 and let $W$ be a torsion free $\operatorname{SL}(2,\mathbb{Z})$-comodule which is of finite rank as $\mathbb{Z}$-module. We say $W$ has a good filtration if there is a finite filtration of comodules $F_{i}\subseteq W$ with
@@
By Corollary 2.9 the following holds: For any pair of integers $1\leq n\leq m$ it follows the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ has a good filtration. We may ask for a classification of torsion free comodules on $\operatorname{SL}(2,\mathbb{Z})$ having a good filtration.
@@
Let $G:=\operatorname{SL}(2,\mathbb{Z})$ and let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ be the standard $G$-comodule from Definition 2.1. Define for any integer $m\in\mathbb{Z}$ a map $f_{m}:V\rightarrow V$ by $f_{m}(v):=mv$. Let $I:=(m)\subseteq\mathbb{Z}$ be the ideal generated by $m$ and assume $I\neq(0),(1)$. Consider the following exact sequence pf $\mathbb{Z}$-modules:
@@
Let $R:=A[x_{ij}]/(D-1)$ with $\operatorname{SL}(2,A):=\operatorname{Spec}(R)$ and let $T:R\rightarrow R$ be the following map:
@@
The transpose of a comodule on $\operatorname{SL}(2,\mathbb{Z})$.
As an example: If $G:=\operatorname{SL}(2,\mathbb{Z})$ it follow any comodule $W$ that is finitely generated as $\mathbb{Z}$-module, is the direct sum (a direct sum of comodules) $W\cong E(W)\oplus T(W)$ with $E(W)\cong\mathbb{Z}^{n}$ a free rank $n$ module on $\mathbb{Z}$. The module $E(W)$ is the torsion free part of $W$. We may for such modules always define the transpose comodule $W^{*}\cong E(W)^{*}$ since $T(W)^{*}=(0)$ is zero. Hence for any comodule $W$ on $\operatorname{SL}(2,\mathbb{Z})$ that is finitely generated as $\mathbb{Z}$-module we may always define the transpose comodule $W^{*}$ using Proposition 4.12.
@@
Let $k$ be the field of rational numbers and let $\operatorname{SL}(2,k)$ be the pull back of $\operatorname{SL}(2,\mathbb{Z})$ to $k$. There is by [11] an isomorphism
$\eta:\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))\cong\operatorname{K}(\operatorname{SL}(2,k))$
@@
The group $\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$ is generated by the classes $[\operatorname{S}^{d}(V)]$ where $d\geq 0$ is an integer and $(V,\Delta)$ is the standard comodule. There is an equality of classes
@@
Let $x\in\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$ be any class. It follows $\eta(x)=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V\otimes k)]$ for some integers $d_{i}\geq 0$ and $n_{i}\in\mathbb{Z}$. It follows $y:=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V)]\in\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$ and since $\eta$ is injective and $\eta(x-y)=0$, it follows $x=y$. The second statement follows using a simlar argument since there is an isomorphism
@@
Note: Lemma 3.15 and Corollary 2.9 suggests that many torsion free comodules on $\operatorname{SL}(2,\mathbb{Z})$ have a good filtration. Since the group scheme $\operatorname{SL}(2,A)$ fails to be completely reducible for a general Dedekind domain, the existence of a good filtration for a large class of torsion free comodules may have interesting consequences for this study. In the following we give more nontrivial examples of torsion free comodules with a good filtration.
@@
Let $(V,\Delta)$ be the standard $\operatorname{SL}(2,\mathbb{Z})$-comodule and let $W_{1},W_{2}$ be $\operatorname{SL}(2,\mathbb{Z})$-comodules on the form $\operatorname{S}^{n}(V),\operatorname{S}^{m}(V^{*})$ or $(\operatorname{S}^{k}(V)^{*},\Delta^{\tilde{T}}_{k})$ for integers $n,m,k\geq 0$. It follows the tensor power $W_{1}\otimes W_{2}$ has a good filtration.
@@
There is an isomorphism of comodules $V\cong V^{*}$ hence $\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V^{*})$ are isomorphic as comodules. If $v_{i}:=e_{1}^{n-i}e_{2}^{i},x_{i}:=v_{i}^{*}$ are bases for $\operatorname{S}^{n}(V)$ and $*\operatorname{S}^{n}(V)^{*}$, there is an isomorphism of $\mathbb{Z}$-modules
$*f:\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V)^{*}$
@@
Note: If $k$ is the field of rational numbers and if $V(k):=V\otimes k$ is the standard comodule on $\operatorname{SL}(2,k)$, there is an isomorphism $V(k)\cong V(k)^{*}$ and $\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k)^{*})$. There is a ”canonical” isomorphism of comodules $*\tau:\operatorname{S}^{n}(V(k)^{*})\cong\operatorname{S}^{n}(V(k))^{*}$ defined by
@@
hence we get a ”canonical” isomorphism $*\tau:\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k))^{*}$.
The isomorphism in Theorem 3.17 differs from $\tau$. There is a map similar to $\tau$ defined for $\operatorname{SL}(2,\mathbb{Z})$ but it is no longer an isomorphism of comodules. The isomorphism $\tau$ is constructed in [7], in section B3, page 476.
4. Generalities on duals and left and right comodules on $\operatorname{SL}(2,\mathbb{Z})$
I this section we include some general results on left and right comodules on Hopf algebras. We also give an explicit construction of the dual $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on a Hopf algebra $R$ in the case when $V$ is a free and finite rank module over the base ring $k$. For the group scheme $\operatorname{SL}(n,\mathbb{Z})$ any torsion free finitely generated comodule is free, hence we get a functorial definition of the dual of any torsion free and finitely generated comodule on $\operatorname{SL}(n,\mathbb{Z})$. Some of the proofs in the section are left to the reader as an exercise. They are standard proofs involving commutative diagrams and a bit of homological algebra.
We use the construction of the dual to give an explicit and elementary construction of a torsion free comodule $W$ on $\operatorname{SL}(2,\mathbb{Z})^{op}$ where $W$ and the dual $W^{*}$ have the same weight space decomposition for the induced action of $\operatorname{Lie}(\operatorname{SL}(2,\mathbb{Z})^{op})$, but where $W$ and $W^{*}$ are non isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z})^{op}$. Hence a torsion free comodule is not determined by its weight space decomposition as is the case over a field $k$ of characteristic zero.
@@ 4. Generalities on duals and left and right comodules on $\operatorname{SL}(2,\mathbb{Z})$
Note: The category $\underline{A-alg}$ of all commutative $A$-algebras is not a small category, hence if you have issues with non standard set theory, you may prefer the approach using comodules. Else you can use the definition given in the books of of Demazure, Gabriel and Jantzen. The approach in this section gives an explicit construction of the dual comodule $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on $\operatorname{SL}(n,\mathbb{Z})$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module. If you take the ”derivative” of the dual action $\eta^{*}$, you get the dual action of the Lie algebra $\mathfrak{sl}(n,\mathbb{Z})$ on $V^{*}$, but there is a strict inclusion $\operatorname{U}(\mathfrak{sl}(2,\mathbb{Z}))\subsetneq\operatorname{Dist}(\operatorname{SL}(2,\mathbb{Z}))$. You may not use the Lie algebra or the enveloping algebra to study the isomorphism class of the dual comodule $(V^{*},\Delta^{*})$. Two comodules $V,W$ on $\operatorname{SL}(n,\mathbb{Z})$ may be isomorphic as representations of the Lie algebra, but non-isomorphic as comodules.
Let in this section $\operatorname{SL}(2,\mathbb{Z}):=\operatorname{Spec}(R)$ where $(R,\Delta_{R},S,\epsilon)$ is the Hopf algebra introduced in section 1. Let $\sigma:R\otimes_{\mathbb{Z}}R\rightarrow R\otimes_{\mathbb{Z}}R$ be defined by $\sigma(a\otimes b):=b\otimes a$ and let $\tilde{\Delta}_{R}:=\sigma\circ\Delta_{R}$. It follows
@@
By definition $V\otimes B\cong B^{2}$. An element $g\in h_{R}(B)$ is by definition a $2\times 2$ matrix $g:=(a_{ij})$ with $det(g)=1$. Hence we may let $g$ act on $V\otimes B$ using matrix multiplication from the left. One checks $\rho$ defines a natural transformation of functors. The natural transformation $\rho$ is by definition a representation of $\operatorname{SL}(2,\mathbb{Z})$ on the module $V$. We may calculate its corresponding comodule and one finds that the comodule $\Delta_{V}:V\rightarrow V\otimes R$ is the following:
@@
We may view the map $S$ as a group scheme version of the involution $\phi$ from Example 4.6. The map $S$ induce a map of group schemes $\phi_{S}:\operatorname{SL}(2,\mathbb{Z})\rightarrow\operatorname{SL}(2,\mathbb{Z})^{op}$ where $\operatorname{SL}(2,\mathbb{Z})^{op}:=\operatorname{Spec}(R)$ with $\tilde{\Delta}_{R}$ as comultiplication. The map $S$ is an isomorphism of Hopf algebras hence $\phi_{S}$ is an isomorphism of group schemes. We may do a similar construction for any affine group scheme $H$.
@@
The dual of the standard comodule $(V,\Delta_{V})$ on $\operatorname{SL}(2,\mathbb{Z})$.
Let us define the left standard comodule $(V,\Delta^{l}_{V})$ on $\operatorname{SL}(2,\mathbb{Z})$ mentioned in Definition 2.1 as the following map:
@@
Note: Any torsion free $\mathbb{Z}$-module is free hence we may define the dual comodule for any torsion free comodule. Note also that the ”classical dual” from Example 2.3 differs from the dual from Definition 4.12 in the sense that the dual is a comodule on the opposite group scheme $\operatorname{SL}(2,\mathbb{Z})^{op}$. The ”naive dual” is a comodule on $\operatorname{SL}(2,\mathbb{Z})$.
@@
When we compose the functors $D^{LR}_{H}$ and $D^{RL}_{H}$ we get canonical isomorphisms of comodules, hence it makes sense to use these functors to define the dual comodule. The definition may be used to define the dual for any comodule $(V,\Delta_{V})$ on $\operatorname{SL}(n,\mathbb{Z})$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module.
@@
Let $\operatorname{SL}(2,\mathbb{Z}):=\operatorname{Spec}(R)$ and let $\Delta:V\rightarrow R\otimes V$ be the following comodule: Let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ with $B:=\{e_{1},e_{2}\},C:=\{1\otimes e_{1},1\otimes e_{2}\}$ as bases and with the following comodule map:
@@
[…]
@@
Write elements $*w\in\operatorname{S}^{2}(V)^{*}$ as column vectors. It follows

### v6

@@ 3. On the ”complete reduciblilty” property for $\operatorname{SL}_{2,\mathbb{Z}}$
where $g$ is the canonical multiplication map and where $f(v):=zv$ where $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}$. It follows the sequence is an exact sequence of $\operatorname{SL}_{2,A}$-comodules where $A$ is an arbitrary commutative unital ring.
@@
and it follows $a:=\frac{n}{n+1}\in A$. From this it follows $\frac{1}{n+1}\in A$. Hence the projection map $\pi$ does not have a section $s$ if $A=\mathbb{Z}$. Hence the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$ does not satisfy the ”complete reducibility” property. If $(p)\subseteq\mathbb{Z}$ is a prime ideal and if $A:=\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at $(p)$, it follows the sequence
@@
does not split since $p$ is not a unit in $A$. Hence $\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ does not satisfy the ”complete reducibility” property. If $G:=\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ there is always an integer $n$ where the projection map $\pi_{n}$ does not split, hence $G$ does not satisfy the ”complete reducibility” property.
@@
Let $m$ be any positive integer and let $p\in\mathbb{Z}$ be a prime number. The group schemes $\operatorname{SL}_{2,\mathbb{Z}},\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ and $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ do not satifsy the complete reducibility property.
@@
If $k$ is a field of characteristic zero and if $V$ is any finite dimensional $\operatorname{SL}(2,k)$-module, it follows there is a direct sum decomposition $V\cong\oplus_{i\in I}\operatorname{S}^{d_{i}}(V)$ of representations. The symmetric powers $\operatorname{S}^{d}(V)$ are irreducible when $V:=k\{e_{1},e_{2}\}$ is the standard $\operatorname{SL}(2,k)$-module. Proposition 3.7 proves this property does not hold for the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$. The symmetric power of the standard module $\operatorname{S}^{d}(V)$ is not irreducible and there is no decomposition of the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ into a direct sum of irreducible modules. By the above example it follows $V\otimes\operatorname{S}^{n}(V)$ is not isomorphic to the direct sum $\operatorname{S}^{n-1}(V)\oplus\operatorname{S}^{n+1}(V)$ as comodules. Hence the ”classification” of $\operatorname{SL}_{2,\mathbb{Z}}$-modules of finite rank over $\mathbb{Z}$ is more complicated than the case of $\operatorname{SL}(2,k)$.
@@
Let $V$ be the standard comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ from Definition 2.1 and let $W$ be a torsion free $\operatorname{SL}_{2,\mathbb{Z}}$-comodule which is of finite rank as $\mathbb{Z}$-module. We say $W$ has a good filtration if there is a finite filtration of comodules $F_{i}\subseteq W$ with
@@
By Corollary 2.9 the following holds: For any pair of integers $1\leq n\leq m$ it follows the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ has a good filtration. We may ask for a classification of torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ having a good filtration.
@@
Let $G:=\operatorname{SL}_{2,\mathbb{Z}}$ and let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ be the standard $G$-comodule from Definition 2.1. Define for any integer $m\in\mathbb{Z}$ a map $f_{m}:V\rightarrow V$ by $f_{m}(v):=mv$. Let $I:=(m)\subseteq\mathbb{Z}$ be the ideal generated by $m$ and assume $I\neq(0),(1)$. Consider the following exact sequence pf $\mathbb{Z}$-modules:
@@
Let $R:=A[x_{ij}]/(D-1)$ with $\operatorname{SL}_{2,A}:=\operatorname{Spec}(R)$ and let $T:R\rightarrow R$ be the following map:
@@
The transpose of a comodule on $\operatorname{SL}_{2,\mathbb{Z}}$.
As an example: If $G:=\operatorname{SL}_{2,\mathbb{Z}}$ it follow any comodule $W$ that is finitely generated as $\mathbb{Z}$-module, is the direct sum (a direct sum of comodules) $W\cong E(W)\oplus T(W)$ with $E(W)\cong\mathbb{Z}^{n}$ a free rank $n$ module on $\mathbb{Z}$. The module $E(W)$ is the torsion free part of $W$. We may for such modules always define the transpose comodule $W^{*}\cong E(W)^{*}$ since $T(W)^{*}=(0)$ is zero. Hence for any comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ that is finitely generated as $\mathbb{Z}$-module we may always define the transpose comodule $W^{*}$ using Proposition 4.12.
@@
Let $k$ be the field of rational numbers and let $\operatorname{SL}(2,k)$ be the pull back of $\operatorname{SL}_{2,\mathbb{Z}}$ to $k$. There is by [11] an isomorphism
$\eta:\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})\cong\operatorname{K}(\operatorname{SL}(2,k))$
@@
The group $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ is generated by the classes $[\operatorname{S}^{d}(V)]$ where $d\geq 0$ is an integer and $(V,\Delta)$ is the standard comodule. There is an equality of classes
@@
Let $x\in\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ be any class. It follows $\eta(x)=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V\otimes k)]$ for some integers $d_{i}\geq 0$ and $n_{i}\in\mathbb{Z}$. It follows $y:=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V)]\in\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ and since $\eta$ is injective and $\eta(x-y)=0$, it follows $x=y$. The second statement follows using a simlar argument since there is an isomorphism
@@
Note: Lemma 3.15 and Corollary 2.9 suggests that many torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ have a good filtration. Since the group scheme $\operatorname{SL}_{2,A}$ fails to be completely reducible for a general Dedekind domain, the existence of a good filtration for a large class of torsion free comodules may have interesting consequences for this study. In the following we give more nontrivial examples of torsion free comodules with a good filtration.
@@
Let $(V,\Delta)$ be the standard $\operatorname{SL}_{2,\mathbb{Z}}$-comodule and let $W_{1},W_{2}$ be $\operatorname{SL}_{2,\mathbb{Z}}$-comodules on the form $\operatorname{S}^{n}(V),\operatorname{S}^{m}(V^{*})$ or $(\operatorname{S}^{k}(V)^{*},\Delta^{\tilde{T}}_{k})$ for integers $n,m,k\geq 0$. It follows the tensor power $W_{1}\otimes W_{2}$ has a good filtration.
@@
There is an isomorphism of comodules $V\cong V^{*}$ hence $\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V^{*})$ are isomorphic as comodules. If $v_{i}:=e_{1}^{n-i}e_{2}^{i},x_{i}:=v_{i}^{*}$ are bases for $\operatorname{S}^{n}(V)$ and $\operatorname{S}^{n}(V)^{*}$, there is an isomorphism of $\mathbb{Z}$-modules
$f:\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V)^{*}$
@@
Note: If $k$ is the field of rational numbers and if $V(k):=V\otimes k$ is the standard comodule on $\operatorname{SL}(2,k)$, there is an isomorphism $V(k)\cong V(k)^{*}$ and $\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k)^{*})$. There is a ”canonical” isomorphism of comodules $\tau:\operatorname{S}^{n}(V(k)^{*})\cong\operatorname{S}^{n}(V(k))^{*}$ defined by
@@
hence we get a ”canonical” isomorphism $\tau:\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k))^{*}$.
The isomorphism in Theorem 3.17 differs from $\tau$. There is a map similar to $\tau$ defined for $\operatorname{SL}_{2,\mathbb{Z}}$ but it is no longer an isomorphism of comodules. The isomorphism $\tau$ is constructed in [7], in section B3, page 476.
4. Generalities on duals and left and right comodules on $\operatorname{SL}_{2,\mathbb{Z}}$
I this section we include some general results on left and right comodules on Hopf algebras. We also give an explicit construction of the dual $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on a Hopf algebra $R$ in the case when $V$ is a free and finite rank module over the base ring $k$. For the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$ any torsion free finitely generated comodule is free, hence we get a functorial definition of the dual of any torsion free and finitely generated comodule on $\operatorname{SL}_{2,\mathbb{Z}}$. Some of the proofs in the section are left to the reader as an exercise. They are standard proofs involving commutative diagrams and a bit of homological algebra.
We use the construction of the dual to give an explicit and elementary construction of a torsion free comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$ where $W$ and the dual $W^{*}$ have the same weight space decomposition for the induced action of $\operatorname{Lie}(\operatorname{SL}_{2,\mathbb{Z}}^{op})$, but where $W$ and $W^{*}$ are non isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. Hence a torsion free comodule is not determined by its weight space decomposition as is the case over a field $k$ of characteristic zero.
@@ 4. Generalities on duals and left and right comodules on $\operatorname{SL}_{2,\mathbb{Z}}$
Note: The category $\underline{A-alg}$ of all commutative $A$-algebras is not a small category, hence if you have issues with non standard set theory, you may prefer the approach using comodules. Else you can use the definition given in the books of of Demazure, Gabriel and Jantzen. The approach in this section gives an explicit construction of the dual comodule $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on $\operatorname{SL}_{2,\mathbb{Z}}$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module. If you take the ”derivative” of the dual action $\eta^{*}$, you get the dual action of the Lie algebra $\mathfrak{sl}(n,\mathbb{Z})$ on $V^{*}$, but there is a strict inclusion $\operatorname{U}(\mathfrak{sl}(2,\mathbb{Z}))\subsetneq\operatorname{Dist}(\operatorname{SL}_{2,\mathbb{Z}})$. You may not use the Lie algebra or the enveloping algebra to study the isomorphism class of the dual comodule $(V^{*},\Delta^{*})$. Two comodules $V,W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ may be isomorphic as representations of the Lie algebra, but non-isomorphic as comodules.
Let in this section $\operatorname{SL}_{2,\mathbb{Z}}:=\operatorname{Spec}(R)$ where $(R,\Delta_{R},S,\epsilon)$ is the Hopf algebra introduced in section 1. Let $\sigma:R\otimes_{\mathbb{Z}}R\rightarrow R\otimes_{\mathbb{Z}}R$ be defined by $\sigma(a\otimes b):=b\otimes a$ and let $\tilde{\Delta}_{R}:=\sigma\circ\Delta_{R}$. It follows
@@
By definition $V\otimes B\cong B^{2}$. An element $g\in h_{R}(B)$ is by definition a $2\times 2$ matrix $g:=(a_{ij})$ with $det(g)=1$. Hence we may let $g$ act on $V\otimes B$ using matrix multiplication from the left. One checks $\rho$ defines a natural transformation of functors. The natural transformation $\rho$ is by definition a representation of $\operatorname{SL}_{2,\mathbb{Z}}$ on the module $V$. We may calculate its corresponding comodule and one finds that the comodule $\Delta_{V}:V\rightarrow V\otimes R$ is the following:
@@
We may view the map $S$ as a group scheme version of the involution $\phi$ from Example 4.6. The map $S$ induce a map of group schemes $\phi_{S}:\operatorname{SL}_{2,\mathbb{Z}}\rightarrow\operatorname{SL}_{2,\mathbb{Z}}^{op}$ where $\operatorname{SL}_{2,\mathbb{Z}}^{op}:=\operatorname{Spec}(R)$ with $\tilde{\Delta}_{R}$ as comultiplication. The map $S$ is an isomorphism of Hopf algebras hence $\phi_{S}$ is an isomorphism of group schemes. We may do a similar construction for any affine group scheme $H$.
@@
The dual of the standard comodule $(V,\Delta_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$.
Let us define the left standard comodule $(V,\Delta^{l}_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$ mentioned in Definition 2.1 as the following map:
@@
Note: Any torsion free $\mathbb{Z}$-module is free hence we may define the dual comodule for any torsion free comodule. Note also that the ”classical dual” from Example 2.3 differs from the dual from Definition 4.12 in the sense that the dual is a comodule on the opposite group scheme $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. The ”naive dual” is a comodule on $\operatorname{SL}_{2,\mathbb{Z}}$.
@@
When we compose the functors $D^{LR}_{H}$ and $D^{RL}_{H}$ we get canonical isomorphisms of comodules, hence it makes sense to use these functors to define the dual comodule. The definition may be used to define the dual for any comodule $(V,\Delta_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module.
@@
Let $\operatorname{SL}_{2,\mathbb{Z}}:=\operatorname{Spec}(R)$ and let $\Delta:V\rightarrow R\otimes V$ be the following comodule: Let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ with $B:=\{e_{1},e_{2}\},C:=\{1\otimes e_{1},1\otimes e_{2}\}$ as bases and with the following comodule map:
@@
as follows. Let $\operatorname{S}^{2}(V)^{*}\cong\mathbb{Z}\{(e_{1}^{2})^{*},(e_{1}e_{2})^{*},(e_{2}^{2})^{*}\}$ with $B_{3}:=\{(e_{1}^{2})^{*},(e_{1}e_{2})^{*},(e_{2}^{2})^{*}\}$ and let $C_{3}:=\{(e_{1}^{2})^{*}\otimes 1,(e_{1}e_{2})^{*}\otimes 1,(e_{2}^{2})^{*}\otimes 1\}$. Define the following matrix:
@@
Write elements $w\in\operatorname{S}^{2}(V)^{*}$ as column vectors. It follows